### 26 Dec 2023 v1

# Towards Flexible, Scalable, and Adaptive Multi-Modal Conditioned Face Synthesis

Abstract
Recent progress in multi-modal conditioned face synthesis has enabled the creation of visually striking and accurately aligned facial images. Yet, current methods still face issues with scalability, limited flexibility, and a one-size-fits-all approach to control strength, not accounting for the differing levels of conditional entropy, a measure of unpredictability in data given some condition, across modalities. To address these challenges, we introduce a novel uni-modal training approach with modal surrogates, coupled with an entropy-aware modal-adaptive modulation, to support flexible, scalable, and scalable multi-modal conditioned face synthesis network. Our uni-modal training with modal surrogate that only leverage uni-modal data, use modal surrogate to decorate condition with modal-specific characteristic and serve as linker for inter-modal collaboration , fully learns each modality control in face synthesis process as well as inter-modal collaboration. The entropy-aware modal-adaptive modulation finely adjust diffusion noise according to modal-specific characteristics and given conditions, enabling well-informed step along denoising trajectory and ultimately leading to synthesis results of high fidelity and quality. Our framework improves multi-modal face synthesis under various conditions, surpassing current methods in image quality and fidelity, as demonstrated by our thorough experimental results.
Figure 1: Our method’s versatile synthesis capabilities, demonstrating high-fidelity facial image generation from a flexible combination of modalities including mask, text, sketch, lighting, expression, pose, and low-resolution images. Remarkably, these diverse face synthesis tasks are achieved within a single sampling process of a unified diffusion U-Net, demonstrating the method’s efficiency and the seamless integration of multi-modal information.
1 Introduction
Diffusion models have made remarkable achievements on a wide range of synthesis tasks in various domains e.g. image  [27, 16], audio [24, 17] ,video [1, 6] and motion [38, 35]. Within the sphere of image synthesis, the focus has increasingly shifted towards more controllable synthesis under multi-modal conditions [37, 21, 22, 20, 10]. These models can generate images that are not only visually compelling but also align well with given multimodal conditions. The emerging developments in this field are setting the stage for more dynamic and user-centric image synthesis, broadening its practicality and impact in real-world applications.
One approach to achieving multi-modal conditioned synthesis, as explored in recent studies [22, 20], involves first developing uni-modal condition synthesis diffusion models and then averaging the noises generated by each modal branch, as shown in Figure 2 (a). This technique allows for flexible multi-modal conditions synthesis, enabling image generation under any combination of modal inputs. However, a notable limitation is its limited scalability. Each modality requires a distinct synthesis network, and the inference complexity increases linearly with the addition of each new modality. Building upon these foundations, Collaborative Diffusion [13] extends the work of [20, 22] by introducing a meta network to fuse the noise from each uni-modal branch, but require independently annotated multi-modal data for training. Another alternative methodology in the realm of multi-modal conditioned synthesis involves incorporating additional control mechanisms into basic synthesis models, as shown in Figure 2 (b). Notable examples include ControlNet [37] and T2I-adapter [21], which introduce layout control into text-conditioned image synthesis frameworks. These innovations enable image synthesis under the combined modal conditions of layout and text. However, this line of methods encounters limitations in terms of flexibility. They require specific tuning for each unique combination of modalities and require multi-modal annotated data. In addition, these methods often neglect the inherent differences in conditional entropy, a measure of unpredictability in data given some condition, associated with conditions from various modalities. As a result, they tend to assign uniform control strength across all modalities, disregarding their unique informational characteristics. In multi-modal synthesis scenarios, applying a fixed control strength uniformly across all modalities is suboptimal. This is particularly evident where conditions of high conditional entropy, under low control strength, fail to exert adequate influence on the image synthesis process. Conversely, an increased control strength might lead to over-fitting in cases involving conditions from modalities with low conditional entropy. Existing multi-modal conditioned synthesis [22, 20, 13] often assigns equal control across modalities, neglecting their distinct conditional entropy. As shown in Figure  5 where higher $w$ indicate higher control strength, a high-entropy modality (text), benefits from stronger control, while masks with lower entropy risk over-fitting.
Figure 2: Core idea comparison between existing multi-modal synthesis approaches and our method. (a) Averaging noises from uni-modal diffusion models for flexible multi-modal synthesis, with scalability challenges due to linearly increased complexity. (b) Incorporation of additional control mechanisms in basic synthesis models for multi-modal synthesis conditioned synthesis, facing flexibility limitations. Each approach contends with the issue of uniform control strength across modalities with varying conditional entropy. (c) Our methods achieve multi-modal conditioned face synthesis within a single synthesis network, under flexible combination of conditions and dynamically adjust noise of diffusion step.
To address the limitations of existing face synthesis frameworks, we introduce a Uni-modal Training approach with Modal Surrogate for enhanced flexibility and scalability, coupled with an entropy-aware modal-adaptive modulation mechanism for responsive adaptation to varying modal conditions. As shown in Figure 2, we set a distinct modal surrogate for each modality, designed to function as both a condition decorator for its respective modality and an inter-modal linker to facilitate collaboration between different modalities. During training, we only use uni-modal annotated data. The surrogate of the active modality is merged with the given condition, along with the surrogates of other modalities, to guide the de-noise process of a single diffusion U-Net. During this process, both the surrogate of the active modality and those of the other modalities are updated. The active modality’s surrogate acquires modality-specific information, while the others learn to interact with the active modality. This cycle is repeated for all considered modalities during training, ensuring each surrogate captures modality-specific context as a complement to the given conditions and serves as an interface for multi-modal conditioned face synthesis. Our synthesis network is highly flexible and scalable, capable of generating facial images under a variety of modal combinations within a single sampling process of a single diffusion model.
Given the disparate conditional entropy inherent to different modalities, we further develop an entropy-aware modal-adaptive modulation. This mechanism thoughtfully adjusts the noise levels, allowing the network to adapt its de-noising strategy to the distinct characteristics of each modality. Consequently, this adaptive approach ensures that each modality effectively influences the image generation process and the synthesis of facial images is faithful to all the given conditions.
Using our novel approach of uni-modal training with modal surrogates, complemented by an entropy-aware, modal-adaptive modulation mechanism, our framework excels in generating high-fidelity facial images across multi-modal conditions. We extend beyond the standard modalities of masks, text and sketch ever considered in existing methods [13, 22, 20, 34, 37, 21], incorporating 3D face models and low-resolution images to enrich the multi-modal synthesis applications. This allows for a wide range of applications, e.g. uni-modal condition face synthesis including facial image super-resolution, lighting and sketch conditioned face synthesis, mask, sketch, text and lighting conditioned guided face generation, and so on, as shown in Figure  1.
Our contributions are summarized as following:
  • We devise an uni-modal training framework that assigns a unique modal surrogate to each modality, greatly enhancing the face synthesis process in terms of flexibility and scalability. These modal surrogates serve dual purposes: acting as condition decorators for their respective modalities and as inter-modal linkers, enabling our network to efficiently process a wide array of modal inputs within a unified diffusion model.
  • We propose an entropy-aware modal-adaptive modulation that dynamically adjusts the noise levels based on the conditional entropy of each modality. This allows our system to customize the denoising process, ensuring effective utilization of specific information from each modality and resulting in high-fidelity image synthesis that faithfully adheres to the given multi-modal conditions.
  • Our framework incorporates conditions of more modalities to significantly broaden the scope of our multi-modal synthesis capabilities, supporting a diverse range of face synthesis under uni-modal and complex multi-modal condition. This demonstrates the versatility and creative potential of our approach.
2 Related Work
2.1 Latent Diffusion Model
Recent advancements in image synthesis have been significantly driven by the development of diffusion models, which have shown remarkable success across various domains of generative tasks [27, 16, 4, 9, 5, 35, 38]. Among these, latent diffusion models [27] stand out as a powerful subclass that operates on compressed representations of data. These models map high-dimensional data into a latent space where the diffusion process is applied, leading to efficient synthesis with reduced computational demands. The compressed latent representations retain essential information while filtering out noise, enabling the models to focus on generating coherent structures in the image synthesis process. This approach has opened up new avenues for creating high-quality images that are both diverse and reflective of complex conditional inputs, setting new benchmarks in the field of generative modeling. We build our multi-modal conditioned face synthesis network upon latent diffusion framework.
2.2 Conditioned Face Synthesis
In the evolving landscape of conditioned face synthesis, researchers have explored a range of controls to guide the synthesis process. This includes face generation under condition of sketches [33, 8], semantic masks [23], 3D face models [3, 2, 28, 5], low-resolution images [31], and textual descriptions [26, 30, 36, 19]. Recent methodologies [34, 22, 20, 13, 21, 37, 29, 12] have introduced multi-modal conditions into the face synthesis framework to achieve rich and flexible control than uni-modal condition. One line of methods [13, 22, 20] achieve multi-modal conditioned synthesis by firstly training uni-modal synthesis network in parallel and then fusing noise of each branch at each step.This technique allows for flexible synthesis across modalities. However, it tends to overlook the shared synthesis properties across different modalities, and suffer from an increased sampling complexity proportionate to the number of modalities. Another approach augments existing diffusion models with additional control mechanisms to handle multi-modal inputs [37, 21]. While this approach processes multiple modal inputs within a single network, it lacks flexibility and often requires additional training for new combinations of modalities. A third approach seeks to unify modal inputs by mapping them into a shared feature space, interpolating between various modality representations [29, 34]. However, this can sometimes overlook the inherent differences unique to each modality. Our method integrate the distinct characteristics of each modality and interaction among them within a unified, adaptive synthesis framework to support multi-modal conditioned face synthesis.
3 Method
Figure 3: Results of multi-modal training. The top row are conditions. The synthesis results are presented in the lower two rows. The resulting network can only synthesize pleasing results given all conditions and much of the guidance comes from the low-resolution image. The synthesis network tend to rely on modality of low condition entropy (LR) for synthesis and thus neglect modality with higher condition entropy.
Algorithm 1 Uni-modal Training with Modal Surrogate
1:repeat
2:     $(x_{0},c_{m}),m\sim\mathcal{U}(\{1,\dots,M\})$
3:     $\triangleright$ Sample from uni-modal datasets
4:     $t\sim\mathcal{U}(\{0,\dots,T\}),\epsilon\sim\mathcal{N}(0,\mathcal{I})$
5:     $x_{t}=\sqrt{\bar{\alpha_{t}}}x_{0}+\sqrt{1-\bar{\alpha_{t}}}\epsilon$
6:     $C_{m}=f(e_{1},\dots,\tau(e_{m}+c_{m}),\dots,e_{M})$
7:     $\triangleright$ Fuse condition and all the modal surrogates.
8:     $L=||\epsilon_{\theta_{u},\theta_{w}}(x_{t},C_{m},t)-\epsilon||^{2}$
9:     $\theta_{*}\leftarrow\theta_{*}-\eta\frac{\partial L}{\partial\epsilon_{\theta_{*}}}\frac{\partial\epsilon_{\theta_{*}}}{\partial\theta},*\in\{u,w\}$
10:     $e^{n}\leftarrow e^{n}-\eta\frac{\partial L}{\partial\epsilon_{\theta}}\frac{\partial\epsilon_{\theta}}{\partial C^{m}}\frac{\partial C^{m}}{\partial e^{n}},n\neq m$
11:     $\triangleright$ Updating surrogate of other modality
12:     $e^{m}\leftarrow e^{m}-\eta\frac{\partial L}{\partial\epsilon_{\theta}}\frac{\partial\epsilon_{\theta}}{\partial C^{m}}\frac{\partial C^{m}}{\partial\tau^{m}}\frac{\partial\tau^{m}}{\partial e^{m}}$
13:     $\triangleright$ Updating surrogate of modality $m$
14:until converged
Figure 4: Uni-modal training with modal surrogate and entropy-aware modal-adaptive modulation mechanism. During training, we randomly sample uni-modal data, of which the condition is fused with its modal surrogate to learn modal-specific intrinsic and other modal surrogate to learn inter-modal collaboration. The fused features are sent to the diffusion U-Net to guide the de-noising process of the corrupted input image. The output noise is further modulated according to the condition features and UNet feature to adaptively adjust noise level given the conditions. The training process is provided in Algorithm  1.
Our goal is to synthesize facial image $x$ given conditions from multiple modalities, e.g. text and mask, which can be formally described as:
$P(x|c_{1},\dots,c_{M}),$ (1)
where $M$ is the number of condition modalities. Recent works [22, 20] sample the multi-condition distribution in Equation 1 using stochastic gradient Langvein sampling and extend uni-modal classifier-free guidance for diffusion model [11] into multi-conditioned version:
$\epsilon(x_{t},t,C)=\sum_{m=1}^{M}(w_{m}+1)\epsilon_{\theta_{m}}(x_{t},t,c_{m})-\sum_{m=1}^{M}w_{m}\epsilon_{\theta_{m}}(x_{t},t).$ (2)
As inllustrated in Figure  2 (a), each modality employs a separate synthesis network, ignoring shared aspects of image generation and causing sampling complexity to rise with more modalities. Control strengths $w_{m}$ remain the same across modalities, disregarding their unique entropy. Collaborative Diffusion[13] improves upon this by learning a noise-weighted network but still depends on individual networks and multi-modal data for training.
A concurrent technological trajectory [37, 21] is to achieve multi-modal conditioned synthesis by adding layout control to a pre-trained text-guide image synthesis framework, and can be described as:
$\begin{split}\epsilon(x_{t},t,c_{l},c_{text})=(w+1)\epsilon_{\theta}(x_{t},t,c_{l},c_{text})-w\epsilon_{\theta}(x_{t},t,c_{l}),\end{split}$ (3)
where the $c_{l}$ is layout condition of a single modality and $w$ is actually control strength related only to textual modality. As shown in Figure  2 (b), such methodologies use a shared synthesis network parameter for multiple modalities. To add layout control of another modality, say $c_{3}$, this line of methods requires tedious respective fine-tuning for synthesis conditioned on other combinations like $c_{2}\&c_{3}$.
To alleviate the aforementioned issues, we instead formulate an efficient and flexible multimodal condition synthesis as shown in Figure  2 (c), which can be described as:
$\begin{split}\epsilon(x_{t},t,C)=(w+1)\epsilon_{\theta_{u},\theta_{w}}(x_{t},t,C)-\\ w(x_{t},t,C)\epsilon_{{\theta_{u},\theta_{w}}}(x_{t},t),\end{split}$ (4)
where $C$ represent any combination of modalities, e.g. $c_{1}$, $c_{1}\&c_{2}$ or $c_{1}\&c_{2}\&c_{3}$. $\theta_{u}$ and $\theta_{w}$ represent parameters of diffusion U-Net and weighting network in entropy-aware modal-adaptive modulation, respectively. We dynamically adjust the noise of diffusion net for different condition inputs. Note that we utilize a single model $\theta$ to handle conditions of different modalities with consistent sampling complexity. We describe in detail how we obtain the flexible and efficient synthesis framework with uni-modal training via modal surrogate in Section 3.1. In Section 3.2, we further elaborate our entropy-aware, modal-adaptive modulation mechanism to achieve multi-modal condition synthesis of high fidelity to given multi-modal conditions.
3.1 Uni-modal Training with Modal Surrogate
A straightforward approach for generating facial images under multi-modal conditions involves using multi-modally annotated data to train a synthesis network that aligns with multiple modalities. Such method may prioritize low-entropy modalities like low-resolution images, neglecting other modalities of higher entropy, as shown in figure 3. In multi-modal training, such shortcut issue arises when the network favors modalities with lower condition entropy to make learning process easier. However, this bias towards a specific modality might make the model less versatile across different modalities, not mentioning its requirements for multi-modal data. To fully learn how conditions guide the face synthesis process for each modality, we only utilize the condition of a single modality to train the synthesis network. Following [20, 22], the target in Equation  1 can be decomposed using Bayes’ theorem as
$\begin{split}P(x|c_{1},\dots,c_{M})=p(x)\prod_{m}{\frac{p(x|c_{m})}{p(x)}},\end{split}$ (5)
where uncondition distribution $p(x)$, uni-modal condition distribution $P(x|c_{m})$, and uni-modal conditional distribution $p(x|c_{1},\dots,c_{M})$ are all parameterized by a single network $\theta$ in our framework. To model multi-conditioned distribution $P(x|x_{1},\dots,x_{M})$ , we need to learn unconditional distribution $P(x)$, uni-modal conditional distribution $P(x|c_{m})$ of each modality. A key challenge is to learn inter-modal collaboration to achieve flexible multi-modal condition synthesis. To this end, we set for each modal $m$ a modal surrogate $e_{m}$ which serves as the modal decorator to represent modal-specific features as well as an interface to collaborate with different modalities. The presence of modal surrogate helps to capture the dependencies among modalities, which is critical when these modalities are actually not entirely independent. We rewrite Equation  5 as:
$\begin{split}P(x|c_{1},\dots,c_{M})=\\ P(x|e_{1},\dotsm e_{M})\prod_{m}{\frac{P(x|c_{m}+e_{m},e_{n}(n\neq m))}{P(x|e_{1},\dotsm e_{M})}}.\end{split}$ (6)
Alongside $c_{m}$ and $e_{m}$, $e_{n},n\neq m$ the surrogate of other modalities are also involved in training.
As shown in Figure  4, we first randomly choose a modality $m$ and here we take the modality mask as an example. The condition mask is sent to a mask encoder to extract condition feature $c_{m}$. The modal surrogate $e_{m}$ and condition feature $c_{m}$ are broadcast added together. In this way, $e_{m}$ decorates its control signal with modal-specific context. Other modality surrogates $e_{n},n\neq m$ are also fused with condition features for modality $m$, resulting in fused condition $C_{m}$. Here we instantiate the fuse process by broadcasting $e_{n},n\neq m$ and concatenating them with features for modality $m$ followed by several linear operations. The fused condition $C_{m}$ are sent to the diffusion U-Net to guide the denoising process of corrupted version $x_{t}$ of image $x_{0}$. The parameter of the diffusion U-Net and both the current training modality and other modal surrogates are updated. $e^{m}$ learns to decorate its signal with its modal-specific intrinsic and $e^{n},n\neq m$ learns to interact with the surrogate $e^{m}$ of the activate modality.
During each uni-modal training iteration for a specific modality $m$, the network not only focuses on the characteristics of modality $m$ but also learns how it relates to other modalities through their surrogates $e_{n},n\neq m$. The presence of other modalities’ surrogate enables the network to grasp the contextual relationships that might exist between different modalities. This strategy ensures that the network is not only trained on the specifics of each modality but also understands how different modalities interact, ultimately leading to a model capable of synthesizing outputs from complex multi-modal inputs. By learning cross-modal interactions in this manner, there is no need to design separate training processes for every possible combination of modalities. This approach prevents the network from becoming overly reliant on some modality that might have lower condition entropy, thereby maintaining versatility across different types of data inputs.
Condition
$w=0$
$w=2$
$w=9$
She is in the thirties. No beard.
Condition
$w=0$
$w=2$
$w=9$
$w=20$
Figure 5: Uni-modal synthesis results given different control strength ($w$). To generate facial images of high fidelity and quality, text guided face synthesis require higher control strength than mask.
3.2 Entropy-Aware Modal-Adaptive Modulation
To synthesize facial images of high fidelity given multi-modal conditions, it is essential to consider conditional entropies across modalities. For uni-modal synthesis, one can freely specify the control strength to obtain results of high quality and fidelity. However, in the multi-modal synthesis scenario, a fixed control strength across all modalities may not be appropriate. For some low control strength, the conditions of high condition entropy fail to provide sufficient influence on image synthesis process. Simply increasing control strength may cause the generated images to be over-fit to some condition of low condition entropy. As shown in Figure  5, an appropriate control strength for textual modality is too strong for mask modality. To address this issue, we devise a entropy-aware modal-adaptive modulation where before the output noise is altered by a fixed, user-defined control strength $w$, it is finely adjusted according to given conditions and current denoising state. We use a weighting network that processes the fused features $C_{m}$, the combination of given conditions and modal surrogates, and the U-Net feature $f_{u}$. This network performs average pooling on these features, concatenates the results, and passes them through several linear layers to produce a weighting vector $w$. The modal surrogates encapsulate modality-specific priors. When combined with the given conditions, they offer crucial guidance on noise adjustment. The inclusion of U-Net features supplies real-time insights $f_{u}$ into the denoising trajectory. In addition to predicting a single base noise map based on U-Net feature $f_{u}$ we predict multiple ($K$) noise maps, each potentially capturing different aspects of the features that could be added to the image at the current step of the diffusion process. We finally obtain the output noise as:
$\begin{split}n=\frac{1}{K}\sum_{k}(w_{k}(n_{k}-n_{b})+n_{b}).\end{split}$ (7)
We start with a base noise pattern $n_{b}$ and introduce adjustment from other noise maps $n_{k}$ of which the influence is scaled by $w_{k}$ according to the specific requirements of the multimodal conditioned image synthesis process. The noise is finely modulated according to modal-specific information, given conditions and current state, enabling well-informed and strategic decisions in each step of de-noising and ultimately leading to synthesis result of high quality and high fidelity to multi-modal conditions.
TediGAN [34]
Compose [20, 22]
CoDiff [13]
T2I-Adapter [21]
ControlNet [37]
This woman looks extremely young. There is no beard on her face
He is in his middle age. The face is covered with beard of medium length.
TediGAN [34]
Compose [20, 22]
CoDiff [13]
T2I-Adapter [21]
ControlNet [37]
Ours
Figure 6: Our methods generate face images of higher quality and align better with the given face parsing map and textual description, compared with existing competitive multi-modal condition synthesis methods.
Method Training Data Type FID $\downarrow$ Text(%) $\uparrow$ Mask (%) $\uparrow$
TediGAN [34] Uni-modal 116.04 24.38 86.37
Compose [20, 22] Uni-modal 127.39 24.22 77.34
CoDiff [13] Multi-modal 122.51 24.37 85.94
ControlNet [37] Multi-modal 136.41 25.70 85.44
T2I-Adapter [21] Multi-modal 139.82 26.11 79.84
Ours Uni-modal 103.14 24.70 90.16
Table 1: Comparison with existing methods on multi-modal conditioned (mask + text) face synthesis.
4 Experiments
4.1 Experimental Setup
Datasets. We conduct experiments on the Celeb-HQ dataset [15] with 30k high-quality face images. We utilized mask from [18] and textual descriptions from [14]. We compared our method, focusing on parsing mask and text modalities, for fair comparison with existing multi-modal methods following [13]. Our approach allowed to efficiently include additional modalities such as sketches [34], 3D face models [7], and low-resolution images [32], enriching our range of synthesis tasks.
Figure 7: Comparison of synthesis results across training methodologies. Pink arrows indicate image generation with mask condition, while blue arrows represent text involvement. Our uni-modal training with modal surrogate enables a flexible and scalable face synthesis framework. Entropy-aware Modality-Adaptive Modulation further enhances the fidelity to given conditions.
Figure 8: Quantitative and qualitative comparison of synthesis results of our full methods and ours without Entropy-aware Modal-adaptive Modulation (EAM) under different control strength. As control strength increases, the fidelity of synthesis results increases and our full methods achieve consistently superior results than ours without EAM.
Implementation details. We implement our method based on Latent Diffusion [27], where the images $(3\times H\times W)$are first mapped to latent features $(3\times\frac{H}{8}\times\frac{W}{8})$. The modal encoder of the low-resolution image is the same as the first stage encoder of LDM [27]. The encoder of text and low-resolution image is fixed during training, while the encoders of other modalities are updated in training. The modal surrogate for text is initialized as the textual encoder feature of input human face and others are randomly initialized.
Evaluation metrics. We assess our method’s performance using Frechet Inception Distance (FID) for image quality and diversity, text matching accuracy through CLIP’s vision-text space cosine similarity (measuring alignment between generated images and text descriptions), and mask accuracy, evaluated by comparing parsing results of generated images against ground-truth semantic maps.
4.2 Comparison Analysis
Compared methods. We compare our method against several leading approaches: TediGAN [34], Composable [20], Multimodal-diff [22], Collaborative Diffusion [13], ControlNet [37], and T2I-adapter [21].
Quantitative comparisons. In Table  1, we benchmark our method’s FID, Text, and Mask accuracy in multi-modal facial synthesis against leading techniques. TediGAN[34] yield impressive results by optimizing each condition and using a high-resolution generator StyleGAN. ControlNet [37] and T2I-Adpter [21] score well in text accuracy, leveraging advanced text encoders[25] and inheriting capabilities from Stable Diffusion[27]. Compose[22, 20] lags behind due to its simplistic fusion of uni-modal model noise, overlooking crucial modality-specific and inter-modal dynamics. Our method, employing uni-modal training with modal surrogates and entropy-aware adaptive noise modulation, achieves superior results by effectively utilizing and integrating multi-modal controls.
Qualitative comparisons. Figure 6 showcases that our methods produce high-quality facial images, faithfully adhering to specified conditions. While other methods struggle with aligning to details like beards or mask shapes, often resulting in poor quality, our approach excels. Our generated images accurately match text descriptions and mask layouts, demonstrating our method’s proficiency in leveraging each modality’s conditions to create precise, high-fidelity facial images. Please refer to our supplementary materials for more comparison results.
Method Task FID $\downarrow$ Text(%$\uparrow$ Mask (%$\uparrow$
Multi-modal training w/o modal surrogate Mask&Text to Image 124.98 24.28 85.44
Mask to Image - - -
Text to Image - - -
Uncondition to Image - - -
Multi-modal training with modal surrogate Mask&Text to Image 112.25 24.42 89.91
Mask to Image 108.69 - 90.06
Text to Image 195.94 23.18 -
Uncondition to Image 167.21 - -
Uni-modal training with modal surrogate Mask&Text to Image 113.29 24.28 89.72
Mask to Image 109.83 - 88.97
Text to Image 114.07 24.48 -
Uncondition to Image 127.26 - -
Uni-modal training with modal surrogate Noise Reweighting Mask&Text to Image 103.14 24.70 90.16
Mask to Image 112.91 - 90.31
Text to Image 110.32 24.78 -
Uncondition to Image 129.52 - -
Table 2: Ablation results of uni-modal training with modal surrogates and entropy-aware modal-adaptive modulation.
4.3 Ablation Analysis
We first validate our uni-modal training with modal surrogates for facilitating flexible and scalable face synthesis. Next, we examine the effectiveness of our entropy-aware modal-adaptive modulation in fully leveraging the control of varing multi-modal conditions.
Figure 9: Demonstrating synthesis capability under the combination of any modalities, including uni-modal conditioned synthesis, pose & text conditioned synthesis, and text & sketch & lighting & mask conditioned synthesis, etc.
Uni-modal training with modal surrogate. Our baseline is to directly train a multi-modal synthesis network using multi-modal data. As shown in the top row of Table  2 and Figure 7 (a), this methods effectively generates images combining mask and text inputs. However, it’s limited to specific mask-text combinations and doesn’t support unconditional or single-modality generation. Enhancing our initial approach, we integrated modal surrogates, enabling the network to combine condition with its corresponding modal surrogate for image generation. This method, shown in the second top-down row of Table  2 and Figure  7 (b), supports not only multi-modal synthesis but also uni-modal or unconditional generation, thanks to the modal surrogate’s representative capabilities. It is noticed a marked performance gap between text-to-image and mask-to-image synthesis: the FID text-to-image was 195.94, much higher than FID for mask-to-image 108.69. This disparity is largely due to the lower conditional entropy of masks, which offer clearer visual cues, lead the network overly rely on mask and thus neglect text. Figure  7 (b) shows that visual results of text-to-image is much worse than mask-to-image. To address this, we designed a uni-modal training mechanism with modal surrogates, yielding impressive results in multi-modal, uni-modal, and unconditional synthesis as shown in the third row from the top of Table  2. This approach leverages information from each modality and uses modal surrogates to learn inter-modal interactions, requiring only uni-modal data.
Entropy-aware modal-adaptive modulation. Enhancing our uni-modal training with modal surrogates, we integrated entropy-aware modal-adaptive modulation to adjust noise before it is altered by a fixed control strength. As illustrated in Table  2 and Figure  8, this method consistently outperforms others in multi-modal, uni-modal, and unconditional synthesis, especially in terms of fidelity to mask and text, while retaining high quality. Our entropy-aware approach ensures that all multi-modal conditions are given due consideration, leading to higher-fidelity image synthesis.
4.4 Extension to More Modalities
Our method’s notable flexibility and scalability enables effective learning of facial image generation from uni-modal training with modal surrogates. Additionally, our entropy-aware, modality-adaptive modulation fine-tunes denoising based on each modality’s specifics, ensuring harmonious multi-modality contributions. This versatility has led to the seamless integration of various controls such as sketches, low-resolution images, and 3D face models, enhancing our network’s capability for diverse facial synthesis tasks.
Diverse uni-modal synthesis. Figure  9 (a) demonstrates our model’s adeptness in uni-modal synthesis, consistently producing high-quality facial images with great fidelity to each modality. Whether it’s face shape from mask, an expression from a 3D face model, or details in a sketch, our single, versatile model captures these nuances effectively, showcasing its capability to maintain diversity while accurately reflecting each modality’s unique features.
Any combination of modal synthesis. Our method’s integration of various modalities allows for versatile combinations in face synthesis. For example, we can blend mask with lighting (Figure 9 (b)) or create faces based on poses and text (Figure 9 (c)). Our approach aids in text-guided super-resolution (Figure 9 (d)), and combines sketches with lighting effects (Figure 9 (e)). It can even merge lighting, text, sketch, and mask inputs to produce diverse, high-quality faces (Figure 9 (f)), demonstrating our method’s impressive capability in facial image synthesis.
5 Conclusion
In conclusion, our research presents a significant advancement in the field of multimodal face synthesis, presenting a highly scalable framework that supports face synthesis of high quality and fidelity under flexible combination of condition. Our approach introduces a uni-modal training framework with modal surrogates for each modality that serve as condition decorator for its modality and an inter-linker to facilitate inter-modal collaboration. The entropy-aware modal-adaptive modulation precisely tunes diffusion noise based on modal characteristics and conditions, enhancing the denoising process for superior synthesis quality. Our method significantly broadens multi-modal face synthesis capabilities, supporting a wide range of synthesis tasks from uni-modal to complex multi-modal combinations.